OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81 summation of the quantity negative nine plus six n from n equals zero to fifteen summation of negative fifty four times n from n equals zero to fifteen summation of negative fifty four times n from n equals zero to infinity summation of the quantity negative nine plus six n from n equals zero to infinity

OpenStudy (solomonzelman):

Can you tell me what the difference is?

OpenStudy (anonymous):

+6

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

What is the \(\large\color{blue}{ a_1}\) ?

OpenStudy (anonymous):

-9

OpenStudy (solomonzelman):

Yes, So you would need to find "what number of term is 81 in a sequence" There is a formula, (which you have already seen most likely). \(\large\color{blue}{ a_n=a_1+d(n-1)}\) Fill in the numbers, \(\large\color{blue}{ 81=(-9)+6(n-1)}\)

OpenStudy (solomonzelman):

Solution for \(\large\color{blue}{ n}\) will give you the the answer, for which(th) term in a sequence is 81.

OpenStudy (anonymous):

16?

OpenStudy (solomonzelman):

yes, I will go over just in case. \(\large\color{blue}{ 81=(-9)+6(n-1)}\) \(\large\color{blue}{ 90=6(n-1)}\) \(\large\color{blue}{ 15=n-1}\) \(\large\color{blue}{ 16=n}\)

OpenStudy (solomonzelman):

So you know that \(\large\color{blue}{ 81}\) is the \(\large\color{blue}{ a_{16}}\)

OpenStudy (solomonzelman):

You have: \(\large\color{blue}{ a_{1}=-9}\) \(\large\color{blue}{ d=6}\) \(\large\color{blue}{ 81}\) is the \(\large\color{blue}{ a_{16}}\)

OpenStudy (solomonzelman):

You want to write the sum for this pattern assuming (where the dots in the questions are, i.e. +3, + 9 ... + 81 ) THAT the pattern is going on (that you add +6 each time).

OpenStudy (solomonzelman):

You are starting from an index (on the bottom of the \(\large\color{blue}{ \Sigma}\) ) of n=0.

OpenStudy (solomonzelman):

(as your options say.

OpenStudy (solomonzelman):

\(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ ? } ~~?}\)

OpenStudy (anonymous):

i think the # on top is 15

OpenStudy (solomonzelman):

Yes, correct.

OpenStudy (solomonzelman):

Because if the first term is given when n=0 (since that is where the index starts) then 16th would be given by n=15.

OpenStudy (anonymous):

is the number on the side (-9+6n)

OpenStudy (anonymous):

because -9 is where it starts

OpenStudy (solomonzelman):

\(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ 15 } (6n-9)}\)

OpenStudy (solomonzelman):

yes, that seems correct.

OpenStudy (anonymous):

I think its 6n+9, not -9

OpenStudy (solomonzelman):

Why?

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

you had it arranged the other way

OpenStudy (solomonzelman):

\(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ 15 } (-9+6n)}\) is how you wrote it, thinking about just how the terms go, as, starting from -9. (And this is correct.) I re-wrote it as, \(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ 15 } (6n-9)}\)

OpenStudy (anonymous):

you're correct

OpenStudy (solomonzelman):

yes, just that \(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ 15 } (6n-9)}\) uses less "buttons" than \(\Large\color{royalblue}{ \displaystyle \sum_{ n=0 }^{ 15 } (-9+6n)}\)

OpenStudy (solomonzelman):

it is all same though:)

OpenStudy (solomonzelman):

just as saying: y=xm-b and y=-b+xm

OpenStudy (solomonzelman):

but very well done, so all we did was, that we found the differences and the first term. Then found which(th) number is the last term in the sequence. After this we wrote the summation.

OpenStudy (anonymous):

I know, i wasn't paying attention

OpenStudy (solomonzelman):

Yw:)

OpenStudy (solomonzelman):

YOu actually came up with a pattern really well!

OpenStudy (anonymous):

you did a good job helping!

OpenStudy (solomonzelman):

That is why I am here.... doing my best:)